# How many qubits does a machine learning problem require?

### v1

@@ 5 Training
In quantum machine learning, models that are expressive are often associated with barren plateaus and thus thought to be untrainable. While this paper focuses on creating a framework for calculating the number of qubits required for a model, we also show a brief demonstration of training with bit-bit encoding on a dataset consisting of first 4 digits of MNIST. We use the model architecture, sub-net initialization, and exact coordinate update optimization techniques from [19].
We create a restricted training set by removing the collisions at each model size. The training batches consist of only the unique samples that occur after collision removal. This is possible as the number of samples grows slower than exponentially with the number of bits as shown in Fig. 4 (a). Interestingly, even when the dataset is fully encoded, the number of unique samples in the encoded dataset is less than the total number of samples in the dataset.
Figure 4: (left) The number of unique samples as a function of the number of bits used to encode the input. The inset shows the same curve with a log scale on the y-axis and compares it with an exponential. (right) Evolution of test accuracy for the first 4 digits of MNIST. $N_{q}$ is the number of qubits. 2 qubits are reserved for reading out the class label, and $N_{q}-2$ qubits are available to load the data sample at the input.
@@ 6 Discussion and Future Directions
In this work, we have examined the universal approximation and resource estimation characteristics of bit-bit encoding for loading classical data into quantum computers. Naively, one may assume that encoding in the basis is useful only when the input data is naturally in the form of binary strings, for example, binary images, combinatorial optimization problems, or molecular orbital occupation vectors. Contrary to this intuition, we have demonstrated that datasets with real-valued entries can also be efficiently represented as bit strings.
@@
Other encoding schemes have been explored with varying degrees of success, but their expressive power is far less predictable, particularly when classical preprocessing and compression are involved. As a result, they lack a systematic framework for resource estimation and provide no reliable way to judge whether the same outcomes could be achieved with fewer qubits. This makes it nearly impossible to meaningfully assess quantum advantage. While such methods may appear to work on small datasets, without a principled understanding of their expressivity, they cannot be expected to scale. Many previous quantum machine learning studies have also leaned on hybrid models, with part of the model parameterized classically and part quantum, which further blurs the line of advantage. In sharp contrast, our model relies exclusively on quantum parameters, making the presence and scope of quantum advantage in learning far clearer.
@@
In the numerical sections of the paper, we studied a variety of encoding schemes, and found that the number of qubits they predict are often very close. Therefore, while we cannot theoretically predict that we have an optimal encoding, it is likely to be close to optimal in an information theoretical sense. Following this, we can be more confident in stating that if a dataset can only fit in a number of qubits that surpasses classical simulation capability of around 50 qubits, it will benefit from quantum advantage within this framework. In particular, it is notable that medium-sized OpenML datasets can be fit within 20 qubits using our method and therefore will not go beyond classical simulation capabilities. This is turn indicates that several datasets that have been studied in the quantum literature so far are not candidates for quantum advantage.
An unexpected finding of our work is that a dataset can sometimes be encoded into fewer number of qubits if it has more dimensions. This is in contrast to our initial expectation that a larger dataset would necessarily require more quantum resources to model. It indicates that the use of the mutual information in the encoding algorithm allows for the efficient extraction of information relevant to the learning task from the original dataset. This phenomenon, to our knowledge, cannot occur with alternative encoding schemes.
For future work, we plan to extend bit-bit encoding to convert more data types into quantum basis states, including sequential, multi-label, regression, and spatial. We also plan to focus on investigating the potential for quantum advantage for datasets from specific fields, such as biology. We will aid this investigation by developing information-theoretic metrics that may support prediction of $Q_{\text{dataset}}$ without iterating through the whole range of qubits, and further validate $Q_{\text{dataset}}$ with training experiments on a variety of datasets.
@@ 7 Software Framework
Figure 5: The number of features (x, log-scale), number of samples (y, log-scale), and number of classes (hue) of each of the $29$ OpenML datasets used in experiments for Section 4.1 and Section 4.2.
Figure 6: The theoretical training and testing accuracy across allocated qubits for the experiment with PCA dimensionality reduction. Each line is the mean for each dataset, with $95\%$ confidence interval error bands.

### v2

@@ 5 Training
In quantum machine learning, models that are expressive are often associated with barren plateaus and thus thought to be difficult to train [mcclean2018barren]. While this paper focuses on creating a framework for calculating the number of qubits required for a dataset, we also show a brief demonstration of training of a quantum model with bit-bit encoding. We use a dataset consisting of first 4 digits of MNIST. We use the model architecture, sub-net initialization, and exact coordinate update optimization techniques from [johri2025bitbitencoding]. We create a restricted training set by removing the collisions at each model size. The training batches consist of only the unique samples that occur after collision removal. This is possible as the number of samples grows slower than exponentially with the number of bits as shown in Figure 5 (left). Interestingly, even when the dataset is fully encoded, the number of unique samples in the encoded dataset is less than the total number of samples in the dataset.
Fig. 5: (left) The number of unique samples as a function of the number of bits used to encode the input. The inset shows the same curve with a log scale on the y-axis and compares it with an exponential. (right) Evolution of test accuracy for the first 4 digits of MNIST. $N_{q}$ is the number of qubits. 2 qubits are reserved for reading out the class label, and $N_{q}-2$ qubits are available to load the data sample at the input.
@@ 6 Discussion and Future Directions
In this paper, we examined the universal approximation and resource estimation characteristics of bit-bit encoding for loading classical data into quantum computers. Naively, one may assume that encoding in the basis is useful only when the input data is naturally in the form of binary strings, for example, binary images, combinatorial optimization problems, or molecular orbital occupation vectors. Contrary to this intuition, we have demonstrated that datasets with real-valued entries can also be efficiently represented as bit strings.
@@
Other encoding schemes have been previously explored with varying degrees of success, but their expressive power is far less predictable, particularly when classical preprocessing and compression are involved. As a result, they lack a systematic framework for resource estimation and provide no reliable way to judge whether the same outcomes could be achieved with fewer qubits. This makes it nearly impossible to meaningfully assess quantum advantage. While such methods may appear to work on small datasets, without a principled understanding of their expressivity, they cannot be expected to scale. Many previous quantum machine learning studies have also leaned on hybrid models, with part of the model parameterized classically and part quantum, which further blurs the line of advantage. In sharp contrast, our model relies exclusively on quantum parameters, making the presence and scope of quantum advantage in learning far clearer.
@@
In the numerical sections of the paper, we applied bit-bit encoding to various real datasets and investigated its performance and dynamics. While we cannot theoretically prove that we have an optimal encoding, if the dimensionality reduction scheme produces independent features, it is close to optimal in an information theoretical sense. As classical simulation hits a wall around $50$ qubits, we can thus state with moderate confidence that if a dataset can requires more than about $50$ qubits to encode, it is a candidate for quantum advantage within this framework. In particular, it is notable that medium-sized benchmark datasets can be fit within $27$ qubits using our method, which implies they will not go beyond classical simulation capabilities. This, in turn, indicates that several datasets that have been studied in the quantum literature so far may not be candidates for quantum advantage. When applying a batched version of bit-bit encoding to a large-scale transcriptomic dataset, we saw the importance of using a dataset with a comprehensive set of samples and informative features. Too many uninformative features will cause bit-bit encoding to estimate a larger $Q_{\text{dataset}}(x)$ than necessary as the PCA and mutual information steps of bit-bit encoding struggle to deal with the noise.
For future work, we plan to extend bit-bit encoding to convert more data types into quantum basis states, including sequential, multi-label, regression, and spatial. We also plan to focus on investigating the potential for quantum advantage for datasets from specific fields, such as biology and chemistry. We will aid this investigation by validating $Q_{\text{dataset}}(x)$ with training experiments that compare the scaling of analogous classical and quantum models.
@@
Finally, in addition to the qubits required for encoding, a full assessment of quantum advantage for a dataset requires analyzing quantities that measure the entangling power, such as operator Schmidt rank, induced multipartite correlation functions, and metrics of non-classicality such as negativity or contextuality, of the classification unitary operator. Upcoming work will compute these quantities to establish a more complete framework for understanding quantum advantage in machine learning.
@@ 7 Software Framework
Fig. 6: The number of features (x, log-scale), number of samples (y, log-scale), and number of classes (hue) of each of the $28$ benchmark datasets used in experiments for Section 4.1 and Section 4.2.
Fig. 7: The theoretical training and testing accuracy across allocated qubits for the experiment with PCA dimensionality reduction. Each line is the mean for each dataset, with $95\%$ confidence interval error bands. The x-axis is cut off at $25$ qubits for increased visibility of individual dataset trends.
Fig. 8: The difference in mean $Q_{\text{dataset}}(1.0)$ distributions of the benchmark datasets when running bit-bit encoding with and without the copula transform.
Fig. 9: The mean $Q_{\text{dataset}}(1.0)$ of the benchmark datasets when running bit-bit encoding with different number of PCA components. The number of components is given by the floor of the number of features multiplied by the component percent.
Fig. 10: The difference between mean $Q_{\text{dataset}}(1.0)$ for each of the benchmark datasets when running bit-bit encoding at $20\%$ component percent and $100\%$ component percent.